Meg 2009 run
Download
1 / 22

MEG 2009 Run - PowerPoint PPT Presentation


  • 142 Views
  • Uploaded on

MEG 2009 Run. Run Coordinator’s View!. DC – HV stability. Review 2009. Successfully finished first MEG Physics Run (Sept. – Dec. 2008). APD - electronics. However – Major Issues to be Investigated/Rectified:

loader
I am the owner, or an agent authorized to act on behalf of the owner, of the copyrighted work described.
capcha
Download Presentation

PowerPoint Slideshow about ' MEG 2009 Run' - kele


An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript
Meg 2009 run

MEG 2009 Run

Run

Coordinator’s

View!

MEG Review February 2010


Review 2009

DC – HV stability

Review 2009

Successfully finished first MEG Physics Run

(Sept. – Dec. 2008)

APD - electronics

  • However – Major Issues to be Investigated/Rectified:

  • DC: high-voltage stability (He-diffusion problem)

  • Calorimeter: LXe light-yield LY (reduced absolute LY  )

  • Calorimeter: PMT gain-drift (reduced gain with >Qbeam)

  • TC: fully functioning fibre detector (new APD readout)

LXe - calibration

MEG Review February 2010


Situation spring 2009
Situation Spring 2009

Post Review February 2009:

Back to “Square One” –

TOTAL Detector DISMANTLED

for Maintenance/Repair /Improvement

Shutdown 2008 / 2009

+

Kottmann Expt. Front-part E5

Scheduled until mid-July!

Calorimeter –

> cooling-power (<LN2)

Modify purification

System – new Getter

(< contamination)

& Liquid Pump (LY)

Dismantling of all DCs

New anode-prints

+ wires +

extensive

tests “aquarium”

& “He-cabin” +…

(HV stability

> )

New APD front-end

electronics

+ control interface

(reduced noise

+ faster -> trig)

MEG Review February 2010


Further planned implementations 2009
Further planned Implementations 2009

  • (I) Upgrading from DRS2 to DRS4

    • differential i/p

    • internal clock & synchronization

    • on-board timing calibration

    • 3.2 GSPS possible XEC-1.6GHz

    • DC  0.5 GHz but >0.7GHz

  • (2) Change of degrader thickness at BTS focus

    • 300m Mylar to 200m Mylar

    • to match 6 % air contamination

    • content in COBRA (previously variable)

    • & centre stopping distribution in target

One main aspect which ran in parallel during the first

half of the year was the ANALYSIS of 2008 DATA!!!!

MEG Review February 2010


Schedule 2009
Schedule 2009

  • Kottmann et al. beam time until mid-July

  • DRS4 installation end-July

  • Parasitic Run (detect. set-up, beam tests e.g. Deg., e+)

  • beg.-Aug.- mid-Sept.

  • e+-test +CEX calibration run mid-Sept. – beg-Oct.

  • Remainder of year

  • - Physics data-taking

  • expectation ~ 2½ months

MEG Review February 2010


Schedule modification
Schedule Modification

  • “Exciting Results” from

  • Lamb-shift Expt.

  • Causeda shift of

  • MEG start up by

  • 3 weeks!!!

  • CONSEQUENCE:

  • In order to enough time

  • To calibrate detector

  • &

  • obtain usable statistics

  • Test beam time reduced

  • to absolute minimum

  • No degrader optimization

  • No positron beam tests

MEG Review February 2010


Organization

Parasitic Run

12 Hr Shifts:

1 DAY SHIFT (Beam Group) 10:00 – 22:00

1 NIGHT SHIFT max. 22:00 – 10:00

1 Shift Leader Nights

Manned by Detector Experts

Basic Run Layout

Parasitic Run

Beam optimization

in parallel with

Debug, Tune &

Calibrate

Full Run Part I

CEX Run + Trigger

Setup + calibrations

Detector monitoring

Full Run

Full Run Part II

Pre-physics data check

Physics Data

(MEG + mixed

triggers)

8 Hr Shifts:

1 DAY SHIFT 07:00 – 15:30

1 EVENING SHIFT 15:00 – 23:30

1 NIGHT SHIFT 23:00 – 07:30

1 Shift Leader + 1 Crew Member

Organization

Similar

structure to 2008

Run Coordinator

2 Shift

Coordinators

 6 weeks tot.

13 Shift Coordinators

 1 weeks/person

Total of 55 persons for 592 shifts (Full Run Only)

  • to allow for flexibility + continuity:

  • Staggered & Overlapping shift system

  • Daily Run Meetings (on-site)

  • Weekly Video Run Meeting (Collaboration-wide)

  • later, weekly Video Physics Analysis Group Meeting

  • Web-based Schedule + Shift list + “On-call” List

MEG Review February 2010


Practical additions monitoring

Access to FE-crates

Access to DC-HV

Access to all

MSCB sub-masters

Practical Additions - Monitoring

  • MIDAS – DAQ control page:

  • web-access to all MSCB sub-masters & nodes

  • web-access to FE crates status, temp, fans, ON/OFF

  • web-access to DC HV-control

  • web-access to Online Analyzer

  • improved Offline analysis jobs for shift crew

  • All towards the long-term GOAL of Remote Shifts….

MEG Review February 2010


Detector synopsis

  • TCs:

  • APD electronics induces too

  • higher noise level in DCs

  •  Switched OFF

  • APD DAQ control problems

  • still Laser temp. control problems

Detector Synopsis

TCs

  • Beam:

  • BTS He cold leak around JT-

  • valve solved temporarily for

  • run

  • Degrader change from

  • 200m to 300m during Run

  • stopping distribution problem

  • Calo:

  • Liquid & Gaseous

  • purification success

  • optimal LY achieved

  • further study PM gain

  • stability

DCs

Beam

  • DCs:

  • HV instability problems

  • solved, fully efficient

  • Resolutions?

Calo.

MEG Review February 2010


Trigger daq synopsis

lcmeg05

lcmeg04

lcmeg03

lcmeg02

lcmeg01

Online

Cluster

Megonxx

Trigger

+ Splitters

Trigger + DAQ Synopsis

  • Trigger:

  • Full Complement of 29 Triggers with pre-scaling used

  • E-resolution improved 20% (7.5%FWHM @ 55MeV)

  •  > thresh.

  • Z-resolution Zrec- ZTC(Qratio) improved ~23% (5.5cm)

  • Direction matching  e with fibresstill missing !!!

  • APD(OFF)  XEC(PMT-index) + TC-bar(index,z)

  • where z from bar charge-ratio

  • Implementation of -trigger during beam=“ON” <TCALIB

  • Trigger#0 TC eff. loss ~25% due broken LVDS transmitter

  • for first-part of the run

  • FE-electronics: mostly NEW!

  • DRS4 teething problems: synchronization + jitter

DRS4 +

part DRS3

  • DAQ:

  • LT~84%, DAQ inter-run time reduced by > 4x

  • DAQ online monitoring & control - many additions

  • Online (backend) 2 TB storage

  • Offline (lcmeg) 64 CPUs + now 150 TB disk

  • “Lazylogger” autocopy Online  Offline

  • factor 2 compression offline

Offline

Cluster

lcmeg

MEG Review February 2010


2009 run conditions

2009 Run Conditions

  • Target Angle: - similar slant angle  as 2008

  • Conventional = (20.6 ± 0.2)°

  • Photogrammetric

  • (outside COBRA) = (20.4 ± 0.2)°

  • Photogrammetric

  • (inside COBRA) = (20.3 ± 0.3)°

2008 <> = (20.5 0.3)º

Target Inclination 2009

 = (20.0 ± 0.3)° provisional

  • Degrader: two settings used during Run

  • 200m & 300m Mylar & 94%He/6%Air

  • Beam Intensity: - on average 10% > IPROT compared 2008  2 main modes tuned

  • “Normal” – same slits as 2008 & “Normal*8” – slits diff., same absolute rate as 2008

Stopping rate for physics Run (300m degrader) at 2.2 mA ~ 2.9·107+s-1

MEG Review February 2010


Degrader question 200 m

t=205 m

l=585 m

Horizontal

Profile

0.8 mm high

Degrader Question 200m

GEM Monte-Carlo 2008 Shows stopping US of target centre by > 50m CH2

for a target thickness of 205 m 4.3% Air

Equivalent to 140 m CH2

Along l or ~ 100 Mylar

300m Degrader

GEM 2008

4.3% Air

Remember 100m of degrader

Equivalent to ~

24% of target thickness ONLY!!!

Therefore should just shift peak of

stopping distribution inside of target

Conclusion:

Remove 100m of Mylar i.e. go to 200m thickness to centre stopping distribution

2009 Physics Run started with 200m Degrader since optimization skipped due to time

Measured

200m

Measured

Vertical

Profile

0.5 mm Aare

Vertex Distribution

Using Michels

MEG Review February 2010


Gem vs data

GEM 2009

6% Air

200

300

strong asymmetry in TC US/DS hit-maps seen???

GEM vs Data

100

300 m Degrader

Example MEG-trigger Data 2009

Run# 59040

200 m Degrader

Example MEG-trigger Data 2009

Run# 53000

  • GEM 2009 –

  • confirms central shift

  • BUT!!!

  • Data features not all

  • reproduced e.g.

  • Asymmetries

  • stopping with no target

Timing problem TC-US

Enhances Asymmetry

For MEG-data

TIC-asymmetry DS/US ~ 4.3

TIC-asymmetry DS/US ~ 1.3

GEM/Data:

*** US-TC

enhancement

MEG Review February 2010


Degrader conclusions
Degrader Conclusions

Motivation:

Originally 2008 degrader of 300 microns changed to 200 microns in 2009 to optimize

stop-distribution to centre of the target for now fixed 6% Air-admixture

amongst other information, this was supported by GEM Monte-Carlo

However – strong asymmetry in TC US/DS hit-maps seen???

Changed back to 2008 situation during Run

Run

Condition

200 Physics Data: 29/10 – 23/11 DAQ-time: 14.0D

300 Physics Data. 26/11 – 23/12 DAQ-time: 21.2D

  • Relative Muon Stopping Statistics 2009/2008:

  • DAQ-time 2009/2008 = 0.732 Dominant factor!

  • Relative time-weighted stopping-rates 2009/2008 = 0.866 (all 300 data = 0.94)

  •  (2009: stop200~56%, stop300~82%) – relative values from Michel data

  • Relative Stopping Statistics: SS2009/SS2008 = 0.634 Dominated by run time!

  • However true relative statistics 2009 much better since spectrometer tracking efficiency

  • & trigger efficiency much better than 2008! factor ~3.5 x 0.634 = ~ Factor 2.2

MEG Review February 2010


Degrader Conclusions cont.

Conclusion: MC does not support “strong leakage” from DS-face for 200m case

detailed check & MC-study underway

Necessary to perform optimization beam test using different degrader thicknesses

Will be performed spring 2010

  • measuring stopping-rate by counting Michel e+ using reconstructed tracks

  • originating from the target volume vs. several degrader thicknesses

  • measure “No-target” stop distribution for above degrader thicknesses to compare

  • with & tune/optimize MC

MEG Review February 2010


Further aspects of 2009 run

Fore-valve

OPEN

CLOSED

isolation vacuum

Further Aspects of 2009 Run

  • BTS: He transfer-line coupling + Joule-Thompson Fore-valve region, cold-leak

  • (meant BTS had to be warmed-up & JT-tower opened + additional pump-stand introduced

  • that could work in B-field environment)

Transfer-line

coupling

FV < 7.5%

He leak Detection

Open: 3·10-7mbls-1

Closed: >7·10-6mbls-1

BTS

JT-Tower

  • Problem solved for Run:

  • New In-coupling transfer line

  • New additional pump-stand

  • for isolation vacuum

FV-OPEN

MEG Review February 2010


Fe electronics
FE-Electronics

  • FE-Electronics:

  • (teething problems with DRS4 implementation took much longer than anticipated

  • involving both software & hardware modifications

  • also the new architecture 4-chips/2-chips -

  • necessitated the total re-doing of the zero-suppression algorithm; physical displacement of

  • cross-associated anode & cathode channels to different modules)

    • DRS4 clock synchronization

    • DRS4 large dispersion in inter –channel, -chip, -board, -crate time-jitter

    • DRS4 spike & ripple suppression

    • DC zero-suppression algorithm

  • Noise problems:

  • (Fourier power spectrum analysis –showed various source frequencies- high frequency

  • range associated with APD electronics resulting in TC fibre system not being used)

    • Environmental noise – seen by DC & TC systems

    • New APD Fe-electronics producer/propagator of high noise-levels on DC

    • readout

  • Details Stefan talk

    Details Flavio talk

    MEG Review February 2010


    Lxe aspects
    LXe Aspects

    Calorimeter:

    2008 Light-yield < expected both for s & s (Q/A)/(Q/A)~ 1.25 expect LP ~ 1.92!!!

    Contamination? new purifier installed Liquid & Gaseous

    & No purification scenarios studied

    • Light-yield proved to be stable @ 1% level during 2009 Run

    • Achieved absolute LY gammas & alphas as expected

    • long-term PMT gain-drift still not understood ~ 6% drop over run

    • monitored extensively by many tools – can be compensated for

    • with HV-adjustment if necessary – problem still being studied

    • gain-variation with beam-loading well understood –new ideas

    GainChronology

    2009 Run

    SKB

    (Q/A)/(Q/A) ~ 1.92 !!!

    SKB = cathode blue sensitivity

    vs. serial no – shows possible link of gain to too much alkali?

    Rel gain-shift

    Serial No

    MEG Review February 2010


    Beam time data

    Run 2009

    93 TB data

    Beam Time/Data

    In view of the extremely tight schedule for starting from “scratch” with the set up in E5,

    also with Lamb-shift extension + numerous detector/electronics etc. problems to be solved … Managed (in real days) 55% of 2008 physics data-taking

    & approximately same number of triggers

    • MEGDATA

    • 200m (Runs# 51824-57184)

    • 5313 Runs a 2k events

    • 10.64 M Triggers

    • Time 14:00.31.17

      300m (Runs# 57582-64482)

    • 5775 Runs a 2k events

    • 11.63 M Triggers

    • Time 21.05.25.07

    Parasitic Run:1st September- 15th September ~ 2 weeks

    Beam Tests/Tuning (~1 week)

    Full Run Part 1:16th Sept. – 30th October ~6 weeks

    CEX 12th October– 24th October (12 days)

    Full Run Part 2:25th October – 22nd December ~8½ weeks

    Physics Data 200m Degrader ~17.5 Days

    Physics Data 300m Degrader ~25.5 Days

    • Normal Physics Data-taking:

    • MEG 12-mixed trigger

    • 6 Hz Trigger Rate, LT~84% 300m Degrader

    • 10 Hz, 75% LT 200m Degrader

    • Daily LED-calibration beam “on-off”

    • 3/week Full-calibration LED beam “on” +LED beam “off”

    • + C-W (Li) + C-W (B) + s

    Total of 93 TB Data

    Taken 2009

    MEG Review February 2010


    Conclusions 2009
    Conclusions 2009

    • Of the original expected ~11½ weeks of physics data-taking (MEG Schedule) minus

    • the 3 weeks extension granted to Lamb-shift expt. We managed to take ~ 8 weeks of

    • MEG data!

    • The major problem encounter during the 2008 Run (DC HV-instability) was solved and the

    • chambers ran with “full efficiency”

    • The LY of the calorimeter also “plateaued” at its expected value allowing /-discrimination

    • with beam “on” enabling continuous monitoring during MEG data-taking – LY stable over

    • extended period of time to ~ 1% level WITHOUT further purification!!!

    • The LXe PMT long-term gain-drift is not understood yet though enough tools available to

    • monitor & sufficient HV-reserve so not problematic. Beam correlated gain variations well

    • studied – strategy for UCN beam structure being further studied.

    • Reduction of the timing-jitter from DRS4 channels on different boards is being studied

    • together with the possibility of running at higher frequency

    • TC fibre-detectors which could not be used in 2009 due to large induced noise on the DC

    • electronics are being modified & expected to be fully functional for 2010

    MEG Review February 2010


    Conclusions 2009 cont
    Conclusions 2009 cont.

    • Fine-tuning of the degrader with beam and a check of the range-straggling in COBRA

    • without a target, for MC comparison/tuning, will be undertaken

    • The postponed test of a monochromatic positron beam which would allow the energy

    • dependent study of our positron spectrometer resolutions as well as the relative

    • acceptances, will be scheduled

    • Many of the outstanding questions concerning detector performance: DC cathode foil

    • aging,TC fibre detector noise suppression, DC-survey position, magnetic field symmetry,

    • chamber resolutions etc. are now being addressed or to be answered soon!

    • – MORE TIME NEEDED!

    • This overview was meant as an introduction to the following “Expert” talks

    MEG Review February 2010


    Finally
    Finally

    A substantial portion of the year was

    ALSO utilized to analyze our

    2008 Physics Data

    In parallel to solving the previous

    outstanding problems.

    The Preprint:

    Is now ready for submission

    for publication!

    MEG Review February 2010


    ad